# Computational Modelling of Thixotropic Multiphase Fluids

### v1

@@ 4.1 Liquid-Liquid Phase Separation (LLPS)
We conduct LLPS simulations in square domains of size $L=10$, in quiescent conditions ($\dot{\gamma}=0$) such that not microstructural breaking is considered. We use a viscosity of the solvent phase $\eta_{s}=1$ and thixotropic transient values for the protein phase. For the protein volume fraction we consider a range $\Phi_{d}=[5\%,12.5\%,25\%,35\%]$, whereas for the capillary velocity $Ca_{v}$, we explore the range $\{0.1,0.5,1,2\}$. To study the effect of the ratio initial-to-final viscosity $\frac{1}{1+\alpha}$ of the protein phase, we use the range $\frac{1}{1+\alpha}=[0.01,0.05,0.1,0.5]$. Finally, to account for the thermal fluctuations we vary the thermal energy in the range $k_{B}T=[0,0.05,0.1]$. To quantify the impact of these variations, we examine both the number $N_{\text{drop}}$ and size $M_{\text{drop}}$ of the droplets formed upon system stabilization, where ”size” refers to the number of particles comprising each droplet. For each type of test, we perform five independent simulations with different random seeds to ensure statistical robustness. To streamline the analysis and comparison between systems, we compute the Probability Density Function (PDF) of the droplets size, and introduce a set of relevant metrics, such as the average molecular weight (AM), the median (MED), the weighted-average molecular weight (WA), and polydispersity index (PDI). These metrics are defined as follows
where $N_{\text{drop}}=\sum N_{i}$ and $M_{i}$ is the size of the $i$-th droplet. A value of $PDI=1$ is an indicative of balanced distribution and homogeneity. In turn, values of $PDI>1$ are linked with high dispersion and greater heterogeneity.
In Figure 5(a), we compile the different Kernel Density Estimation (KDE) curve obtained for the different volume fractions. An snapshot of the final state for two concentration is included, where droplets have different colors for clarity. In B (Figure 13(a) and Figure 13(b)), we provide an example of droplet sizes histograms for $\Phi_{d}=5\%$ and $\Phi_{d}=35\%$, illustrating the central tendency statistics.The results indicate that at the lowest volume fractions ($\Phi_{d}=5\%$) investigated, interfacial tension effects are not sufficient to drive the formation of well-defined droplets. In contrast, at higher volume fractions ($\Phi_{d}=35\%$), the protein phase exhibits a tendency to form regular circular droplets, driven by interfacial tension phenomena. Overall, at lower proteins content lead to greater heterogeneity (broader droplet size), whereas as the volume increases, the droplet size distribution tend exhibit lower dispersions.
(a) Protein phase concentration
(b) Capillary velocity
(c) Constitutive constant $\frac{1}{1+\alpha}$
(d) Thermal fluctuations $k_{B}T$
Figure 5: KDE comparison for the different properties analyses including the PID value.
The capillary velocity ($Ca_{v}$) is a property that, as described by Elbaum-Garfinkle et al. [47] and Mitrea et al. [48], is derived from the relationship between surface tension $\sigma_{0}$ and solvent viscosity $\eta_{s}$, mathematically stated as $Ca_{v}=\sigma_{0}/\eta_{s}$. It exhibits linear behavior in relation to the relaxation time and to the characteristic length of the interacting phases. In Figure 5(b), we apply the same statistical analysis for different values of the capillary velocity, namely $Ca_{v}=[0.1,0.2,1,2]$. To facilitate interpretation, histograms of the cluster size distribution and central tendency statistics are provided for $Ca_{v}=0.1$ and $Ca_{v}=2$ in B (Figure 13(c) and Figure 13(d), respectively). These results reveal that for low $Ca_{v}$ values, clusters exhibit irregular shapes and poorly defined structures, resulting in a broad and heterogeneous size distribution skewed toward smaller clusters. In contrast, higher $Ca_{v}$ values promote the formation of well-defined and uniformly sized clusters, driven by the balance between viscosity and surface tension. Overall, low $Ca_{v}$ values are associated with increased heterogeneity in cluster size, whereas higher $Ca_{v}$ values result in more homogeneous and narrowly distributed cluster populations.
In Figure 5(c), we calculate the KDE curves obtained for different values of the constitutive parameter, expressed as $1/(1+\alpha)$, with values $[0.01,0.1,0.2,0.5]$. Representative histograms and central tendency statistics for $0.011/(1+\alpha)=$ 0.01 and $0.51/(1+\alpha)=$ 0.5 are shown in B (Figure 13(e) and Figure 13(f), respectively). The results indicate that, for higher values of $1/(1+\alpha)$, the clusters exhibit well-defined structures with a more uniform and narrowly distributed size profile. Conversely, lower values of $1/(1+\alpha)$ result in poorly developed structures and broader, less concentrated size distributions. Overall, the data show that as $1/(1+\alpha)$ increases, the system becomes more homogeneous, whereas lower values lead to greater heterogeneity in cluster size.
In Figure 5(d), we compile the KDE curves obtained for different values of the thermal fluctuation parameter, $k_{B}T=[0,0.05,0.1]$, as introduced through the stochastic term in the SDPD equations (2). Representative histograms and central tendency statistics for $k_{B}T=0.05$ and $k_{B}T=0.1$ are shown in Appendix B (13(g) and Figure 13(h), respectively). To allow comparison, we run all simulations for 25 time units, as the system would otherwise converge toward a single, large cluster encompassing the entire dispersed phase. In the absence of thermal fluctuations ($k_{B}T=0$), and under conditions where $Ca$ and $\alpha$ do not significantly affect cluster formation, the system exhibits fully developed, circular structures with relatively small and uniform cluster sizes. As thermal fluctuations increase, clusters become less defined and more variable in size, depending on the time evolution. Higher values of $k_{B}T$ accelerate the merging process, eventually leading to a single dominant cluster. This behavior is reflected in the KDE curves, which show that decreased thermal fluctuations result in higher polydispersity and greater heterogeneity in cluster size before coalescence occurs.
@@ 4.2 Thixotropic flows in channels: Emulsions and suspensions
The behavior of rigid particles suspended in a fluid was studied by Einstein [51], where the expression $\eta_{eff}=\eta(1+B\phi)$ was derived. This indicates that for concentrations less than 1 ($\phi<<1$) the effective viscosity of the particle suspension is directly proportional to the viscosity of the fluid and the volume fraction of rigid particles, with a constant value $B=2.5$. For deformable particles, Taylor [54] proposed:
where $\eta^{\prime}$ is the viscosity of the dispersed phase and $D$ is the system dimension. When $\eta^{\prime}=\eta$, this simplifies to $\eta_{eff}=\eta(1+1.75\phi)$. This model assumes moderate droplet deformability, i.e., low Capillary numbers. We use this expression to validate our multiphase model for emulsions with droplets. We conduct simulations in a square channel of size $L=10$ under simple shear flow $\dot{\gamma}=0.0166$, with particle density $n_{eq}=25$ and no thermal fluctuations. The continuous phase is Newtonian with $\rho=1$, $\eta=15$ and inlet surface tension $\sigma_{0}=2$. Droplets have radius $r_{0}=0.8$ and viscosities $\eta^{\prime}=[15,30]$ covering two viscosity ratios of $\lambda=[1.2]$. This leads to $Re=1.5\times 10^{-3}$ and $Ca=0.1$. We perform a total of eleven simulation with volume fractions in the range $0.30<\phi<$ 0.3 by adjusting the number of droplets. The effective viscosity is calculated by direct measurement of the force on the wall. Figure Figure 6(a) shows the results, including snapshots for $\phi=[0.02,0.14,0.29]$, illustrating droplets deformation. The model captures the increasing trend in viscosity, followed by a slight drop for volumen fraction exceeding 0.2, consistent with prior observations [55, 56, 57, 58].
Once we validate the method for Newtonian cases, we analyze thixotropic flows. Here, the continuous phase has a variable viscosity governed by a thixotropic model, while the dispersed phase retains a constant Newtonian viscosity. We use a channel of length $L=10$, set the continuous phase viscosity to $\eta=1$, and choose thixotropic parameters $\alpha=4$, $\lambda_{0}=100$, and $\beta=1$. These satisfy the condition $\lambda_{0}>L^{2}\rho/\eta_{\infty}$ to guarantee a scale separation between the long microstructural time scale and the short viscous ones [62]. We simulate a simple shear flow ($\dot{\gamma}=0.5$) where a thixotropic continuous phase carries circular dispersed droplets, similar to the Newtonian case. The dispersed phase has a viscosity of $\eta^{\prime}=5$. We test five volume fractions and monitor the emulsion behavior over time. First, we run the system in its Newtonian state until it reaches a steady condition, then switch to the thixotropic model. Results are shown in Figure 6(b). We observe a drop from a peak effective viscosity to a lower steady value as thixotropy develops. Higher concentrations lead to higher effective viscosities.
(a) Droplet suspensions
(b) Thixotropic continuous phase
@@
Figure 6: Emulsion experiments. (a) Validation using Taylor‘s expression for two viscosity ratios $\lambda$ and (b) transient viscosity evolution in a channel with Thixotropic continuous phase for different values of $\phi$
Next, we study the behaviour of emulsions subjected to a Poiseuille flow, starting with the single-phase case. Using a channel of height $L=18$, we apply a body force $g_{x}$ that produces a parabolic velocity profile with a peak value of 6 (in system units). Boundary conditions follow those used in the dynamic validation section. We consider two cases. In the first, we simulate a one-phase flow that begins as Newtonian ($\eta=1$). After reaching steady state, we switch to a thixotropic type with parameters $\alpha=4$, $\lambda_{0}=100$, and $\beta=1$. For comparison, we also run a separate simulation using the maximum viscosity to obtain the limiting velocity profile. Figure 7(a) shows the velocity at $t=100$ (Newtonian state, solid black line) and $t=150$ (thixotropic state, minimum profile). As expected, the velocity decreasesmodel over time as the viscosity increases toward the channel center, reflecting the inverse relationship between viscosity and shear rate in thixotropic fluids.
The second case analyzed has the same objective but is performed for a Poiseuille flow with two phases A and B. We apply the same treatment to a simulation that starts in a Newtonian state and then undergoes transition to the thixotropic model. In this case, phase A maintains its Newtonian behavior throughout the entire simulation, while phase B becomes thixotropic with parameters $\alpha=10$, $\lambda_{0}=100$, and $\beta=1$. As in the previous case, we perform a separate simulation to observe the velocity profile obtained with the maximum viscosity. Figure 7(b) shows the transition between the maximum and minimum velocity profiles. We observe that, as time progresses, the characteristic profile for multiphase flows gradually emerges, as described earlier in the validation section.
(a) One phase
(b) Two phases
Figure 7: Velocity profiles in (a) one and (b) two phases flow. Transition from steady state (both Newtonian fluids) to thixotropic behavior. For the multiphase case only the bottom phase is thixotropic, the upper phase remains Newtonian.
Once we analyze the simple thixotropic cases, we proceed to study emulsions containing droplet suspensions undergoing a Poiseuille flow. For these cases, we use a box of length and height $L=20$ with an array of 25 equidistant and initially undeformed droplets. We consider the continuous phase as Newtonian with viscosity $\eta=5$ and density $\rho=1$, while the dispersed phase exhibits thixotropic behavior with parameters $\eta_{\infty}=5$, $\alpha=4$, $\beta=100$, and $\lambda=100$. We apply the respective body force on the continuous phase to generate the corresponding Poiseuille flow. We analyze two particular cases: one with the formed microstructure, i.e., $f_{0}=1$, and one with the unformed microstructure, $f_{0}=0$. The results appear in the top and bottom rows of Figure 8, respectively. We observe that droplet deformation is smaller in the center of the channel and increases closer to the walls due to the magnitude of the shear rate. Additionally, droplet deformation is smaller in the case with formed microstructure ($f_{0}\approx 1$) since it starts immediately with the highest viscosity and consequently offers a larger hydrodynamic resistance to flow. Conversely, in the case with unformed microstructure ($f_{0}<<1$), the deformation is larger because it starts with the lowest viscosity and evolves over time until reaching its maximum value.
@@
Figure 8: Emulsion with Newtonian continuous phase and thixotropic dispersed phase. Comparison starting with the formed microstructure (top row) versus the unformed microstructure (bottom row).

### v2

@@ 4.1 Liquid-Liquid Phase Separation (LLPS) and aging
We quantify the impact of these variations, by estimating both the number $N_{\text{drop}}$ and size $M_{\text{drop}}$ of the droplets formed up to a final time of $t=25\lambda_{thix}$, where ”size” refers to the radius of the droplet calculated as $R=(\sqrt{N_{part}/d_{eq}})/\pi$. For each type of test, we perform five independent simulations with different random seeds to ensure statistical robustness. To streamline the analysis and comparison between systems, we compute the average molecular weight (AM), the median (MED), the weighted-average molecular weight (WA), and polydispersity index (PDI) (see B).
In Figure 5(a), we summarize the Kernel Density Estimation (KDE) droplet size distributions for systems with varying $\lambda_{\text{thix}}/\tau_{\sigma}$, along with the snapshots of the final state for two different systems (central tendency statistics for $\lambda_{\text{thix}}/\tau_{\sigma}=0.1$ and $\lambda_{\text{thix}}/\tau_{\sigma}=2$ are illustrated in B in Figure 14(c) and Figure 14(d), respectively). These results reveal that for $\lambda_{\text{thix}}/\tau_{\sigma}\leq 0.1$, the droplets exhibit irregular shapes, resulting in a broad and polydisperse size distribution (PID $\sim 1.4$) skewed toward smaller droplets. In contrast, for $\lambda_{\text{thix}}/\tau_{\sigma}\geq 0.5$, the phase separation promotes the formation of well-defined droplets. The increase in $\lambda_{\text{thix}}/\tau_{\sigma}$ beyond $0.5$ seems to not further alter the droplet size distribution, with a PID $\sim 1.1$. This behavior suggests that once the coalescence time scale increases (larger $\lambda_{\text{thix}}/\tau_{\sigma}$) than the characteristic thixotropic restructuring time ($\lambda_{\sigma}<\lambda_{\text{thix}}$), the dynamics of the system are governed primarily by thixotropic kinetics. Thus, further increases in $\lambda_{\text{thix}}/\tau_{\sigma}$ do not accelerate droplet coalescence, since the aging of the protein phase sets the effective upper limit for phase separation dynamics.
The variation of the ratio between the thixotropic and interfacial tension time scales, $\lambda_{\text{thix}}/\tau_{\sigma}$, shows the principal relation between microstructural restructuring and capillary-driven coalescence. When $\lambda_{\text{thix}}/\tau_{\sigma}\leq 0.1$, the interfacial dynamics evolve much slower than the microstructural relaxation, leading to incomplete coalescence events and irregular droplet morphologies. In this regime, extremely polydisperse and kinetically trapped configurations are produced as a result of the rapid interface formation that predominates before the material can reconstruct its internal structure. The two mechanisms operate over similar time scales as the ratio rises approaching $\lambda_{\text{thix}}/\tau_{\sigma}\approx 0.5$, enabling a more balanced evolution in which thixotropic rebuilding gradually stabilizes the droplets while surface tension efficiently reduces interfacial energy. Beyond this threshold ($\lambda_{\text{thix}}/\tau_{\sigma}\gtrsim 1$), the system reaches a quasi-steady regime in which droplet growth and coalescence are limited by the slow microstructural relaxation. In this limit, the thixotropic aging becomes the dominant factor controlling the overall kinetics of phase separation, and further increases in $\lambda_{\text{thix}}/\tau_{\sigma}$ do not significantly alter the droplet size distribution.
In Figure 5(b), we present the variations in the KDE curves for different values of the initial-to-final viscosity ratio $1/(1+\alpha)$. Representative histograms and central tendency statistics for $1/(1+\alpha)=0.01$ and $1/(1+\alpha)=0.5$ are also shown in B (Figure 14(e) and Figure 14(f), respectively). The results indicate that lower initial-to-final viscosity ratios $1/(1+\alpha)\leq 0.01$ lead to poorly developed structures with a larger molecular weight and broader distribution (PID =1.7=1.7). In contrast, as $1/(1+\alpha)$ increases the polydispersity of the systems reduces, favoring narrower size distributions. At the fixed $\lambda_{\text{thix}}=1$ used, a lower initial-to-final viscosity ratio implies that the change in viscosity of the protein phase is significant as the microstructure changes, leading to a very fast reduction in the capillary velocity, and consequently shifting the phase separation time scales towards $\lambda_{\text{thix}}\ll\lambda_{\sigma}$. As a result, the propensity of the system to phase separate decreases rapidly with the system aging, leading to poorly developed polydisperse structures. Conversely, when $1/(1+\alpha)\gtrsim 0.1$, the relative change in viscosity is less pronounced, so that $\lambda_{\text{thix}}/\tau_{\sigma}$ remains sufficiently high during phase separation.
[…]
In the absence of thermal fluctuations, we have that $R^{2}/D_{p}\gg 1$ and the system yields relatively small, well-defined droplets with uniform size distributions, with droplet growth controlled mainly by $\lambda_{\text{thix}}/\tau_{\sigma}$. The resulting distribution remains relatively uniform, but coarsening is incomplete within the simulated time. Decreasing $\tau_{\text{t}}$ (i.e. increasing the $k_{B}T$), enhances diffusive transport, thereby reducing the coarsening time scale $\lambda_{\sigma}\sim R^{2}/D_{p}$. In Figure 5(c), we can observe that the mean size and distribution of the droplets consistently increases with $k_{B}T$. However, the polydispersity index do not change significantly. When $\lambda_{\sigma}$ becomes shorter than $\lambda_{\text{thix}}$, droplets coalesce more rapidly leading to larger droplets, and the dynamics variation in $\lambda_{\text{thix}}/\tau_{\sigma}$ (due to the change in viscosity) do not dominate the phase separation. Thermal fluctuations can dynamically modulate the interplay between $\lambda_{\sigma}$ and $\lambda_{\text{thix}}$. At high $\tau_{\text{t}}$ (low $k_{B}T$) the dynamics are limited by thixotropy, whereas as the coarsening of the phases is favored, the phase separation progresses rapidly enough to control the system’s evolution.
The variation of the thermal time scale $\tau_{\text{t}}$ modulates the balance between diffusive transport, interfacial dynamics, and thixotropic restructuring. For large $\tau_{\text{t}}$ (low $k_{B}T$), phase separation is mainly governed by $\lambda_{\text{thix}}$ and $\tau_{\sigma}$, yielding slow coarsening and uniform droplets. When $\tau_{\text{t}}$ is reduced, enhanced diffusion accelerates droplet coalescence, shortening $\lambda_{\sigma}$ and facilitating more rapid droplet growth. Once $\lambda_{\sigma}$ becomes comparable to or smaller than $\lambda_{\text{thix}}$, the influence of thixotropy weakens and thermal diffusion governs the phase separation. Consequently, $\tau_{\text{t}}$ sets the transition between a regime governed by thixotropy and one controlled by diffusion.
@@ 4.2 Thixotropic flows in channels: Emulsions and suspensions
We now extend the multiphase thixotropic model to describe complex emulsions and suspensions. These systems are of broad relevance since thixotropy has been extensively reported in a several contexts. For example, biological suspensions such as blood and mucus exhibit pronounced time-dependent viscosity, where microstructural rearrangements under flow lead to reversible decreases in viscosity followed by recovery at rest baskurt2003 , lai2009 . Similarly, a wide range of synthetic materials—including concentrated emulsions, colloidal dispersions, and clay-based suspensions—have been shown to display aging, shear-induced breakdown, and subsequent structural rebuilding, all of which give rise to characteristic thixotropic rheology barnes1997 , mewis2009 .
We begin by considering the simplest configuration of a fully separated two-phase system (phases $\phi_{A}$ and $\phi_{B}$) under Poiseuille flow. To avoid spurious artifacts from numerical stabilization, the simulations are initialized with the corresponding stabilized Newtonian velocity profile. In this setup, phase $\phi_{A}$ maintains Newtonian behavior throughout the simulation, while phase $\phi_{B}$ evolves according to the thixotropic model. We employ a square channel with a characteristic length of $L/\Delta x=50$ and a maximum velocity of $U=10$, resulting in a characteristic flow time scale of $\tau_{\text{flow}}=1$. Also we set a sonic time scale $\tau_{\text{c}}=0.2$ and thixotropic time $\lambda_{\text{thix}}=100$ follow the hierarchy $\tau_{\text{c}}\ll\tau_{\text{flow}},\lambda_{\text{thix}}$. We analyze the effects of microstructure destruction to formation rates by setting $\beta=1$, the constitutive parameter $\alpha=4$, and with initially formed ($f_{0}=1$) microstructure. This selection of parameters results in a Reynolds number $Re=1$ and a thixotropic number $Th=1$. For reference, we also perform a simulation in which $\phi_{B}$ remains Newtonian but with its viscosity fixed at the maximum value prescribed by the thixotropic law. As illustrated in Figure 6, the velocity profile in the thixotropic case gradually evolves from the Newtonian baseline toward the limiting profile corresponding to the maximum viscosity. This evolution reflects the interplay of time scales because the convective time $\tau_{\text{flow}}$ is much shorter than the thixotropic relaxation time $\lambda_{\text{thix}}$. The microstructure does not fully rebuild during the initial advection of the fluid, resulting in a transient velocity profile. Over longer times, the microstructure reaches its equilibrium, increases the effective viscosity, and stabilizes the velocity profile. These results demonstrate that the model accurately captures the emergence and stabilization of interfaces in multiphase Poiseuille flow under thixotropic conditions.
Figure 6: Velocity profiles for a two phases flow
Now, we proceed to explore emulsions of i) Newtonian droplets in thixotropic continuous phase, and ii) thixotropic droplets in a Newtonian continuous phase. To avoid the merging of the droplets during flow, we introduce an additional pair-wise soft repulsive force $f_{\text{drop-drop}}=\epsilon_{\text{drop}}(1-r/r_{c})^{2}$, between particles belonging to different droplets, where $r_{c}=4\Delta x$ is the range of the interaction, and $\epsilon_{\text{drop}}$ is the strength of the repulsion that we normalize by a characteristic capillary force scale $\sigma_{0}/r_{c}$, so that $\tilde{\epsilon}_{\text{drop}}=\epsilon_{\text{drop}}r_{c}/\sigma_{0}$. In the present simulations, we set $\epsilon_{\text{drop}}=2.0$, which corresponds to $\tilde{\epsilon}_{\text{drop}}\approx 8$. While this force ensures droplets stability do not introduce any additional artifact in our simulations, as we illustrate in C for fully Newtonian emulsions, where we show that the model consistently reproduces variations of the emulsion’s viscosity with the droplet concentration, as theoretically predicted by Taylor taylor1932viscosity for deformable particles.
First, we consider an emulsion composed of Newtonian droplets ($\eta^{\prime}=5$) suspended in a thixotropic continuous phase. The system is subjected to a simple shear flow ($\dot{\gamma}=0.5$) in a channel of length $L/\Delta x=50$, resulting in a characteristic flow time scale of $\tau_{\text{flow}}=2$. The hierarchy of relevant time scales satisfies $\tau_{\text{c}}\ll\tau_{\text{vis}},\tau_{\sigma}\ll\tau_{\text{flow}},\lambda_{\text{thix}}$, where the sonic, viscous, interfacial, and thixotropic times are set to $\tau_{\text{c}}=0.022$, $\tau_{\text{vis}}=0.04$, $\tau_{\sigma}=0.2$, and $\lambda_{\text{thix}}=100$, respectively. To guarantee a scale separation between the long microstructural time scale and the short viscous ones rossi2022sph , we set the parameters of the thixotropic model such that $\lambda_{\text{thix}}>L^{2}\rho/\eta_{\infty}$. The continuous phase viscosity is set to $\eta_{\infty}=5$, with $\beta=1$, constitutive parameter $\alpha=4$, and an initially formed microstructure ($f_{0}=1$). The radius of the droplets is calculated as $R/\Delta x=(\sqrt{N_{part}/d_{eq}})/\pi=2.2$. The chosen parameters result in a Reynolds number $Re=5$ and a Thixotropic number $Th=0.5$. The systems are initialized as Newtonian emulsion, until the steady state of their velocity profile is reached. Then the thixotropic model is used for the droplet fluid.
We test five volume fractions $\phi$ and monitor the emulsion behavior over time. This volume fraction is calculated as the volume of the dispersed phase ($V_{dp}$) relative to the volume of the continuous phase ($V_{cp}$) i.e. $\phi=V_{dp}/V_{cp}$. In Figure 7, we present the response in the measured viscosity, showing an initial drop from a peak effective viscosity to a lower asymptotic value as the microstructure relaxes from its initially formed state ($f_{0}=1$). This decay reflects the progressive breakdown of the microstructure and the transition of the system toward its thixotropic equilibrium. As expected, higher droplet concentrations lead to higher effective viscosities due to enhanced hydrodynamic interactions. Overall, the plot demonstrates that thixotropy not only reduces or increase (depends of $f_{0}$) the steady-state viscosity but also introduces a transient regime where the effective viscosity evolves dynamically, underlining the role of the thixotropic timescale in determining the apparent rheological response of the system. The observed transient behavior of the effective viscosity can be directly interpreted in terms of the hierarchy of time scales. The droplets are transported along the channel while largely preserving their original shape because the convective time scale $\tau_{\text{flow}}$ is larger than the interfacial time $\tau_{\sigma}$. The longest characteristic time, the thixotropic relaxation time $\lambda_{\text{thix}}$, governs how rapidly the microstructure responds to flow-induced stresses. Given that $\lambda_{\text{thix}}\gg\tau_{\text{flow}}$ and that the system starts from a fully structured state ($f_{0}=1$), the evolution is dominated by microstructure breakdown rather than rebuilding. Consequently, the effective viscosity decreases dynamically as the initially structured droplets progressively adapt to the local flow conditions.
@@
Figure 7: Transient viscosity evolution in a channel with Thixotropic continuous phase for different values of $\phi$
In the case of emulsions of thixotropic droplets suspended on a Newtonian fluid, we simulate a Poiseuille flow in a domain of size $L/\Delta x=100$ containing $25$ initially spherical droplets. In a steady state, the fluid reaches a maximum velocity of $U=1.25$, resulting in a characteristic flow time scale of $\tau_{\text{flow}}=16$. Although the specific value of $\tau_{\text{flow}}$ differs from the previous configuration, the relative ordering of the characteristic times remains unchanged, preserving the hierarchy $\tau_{\text{c}}\ll\tau_{\text{vis}},\tau_{\sigma}\ll\tau_{\text{flow}},\lambda_{\text{thix}}$. The continuous phase has a viscosity $\eta_{s}=5$ and density $\rho=1$, while the dispersed phase is characterized by $\eta_{\infty}=5$, $\alpha=4$ and $\beta=10$. Both, initially fully-formed microstructure ($f_{0}=1$) and unformed ($f_{0}=0$) are considered. As previously calculated, the radius of the droplets is equivalent to $R/\Delta x=2.2$ . Again, we set the parameters of the model to satisfy the condition $\lambda_{\text{thix}}>L^{2}\rho/\eta_{\infty}$ rossi2022sph . This set of parameters yields a Reynolds number of $Re=5$ and a Thixotropic number of $Th=0.625$.
Figure 8 compares the transient evolution of systems starting with a formed microstructure (top row) and an unformed microstructure (bottom row). The droplet deformation depends on the local shear rate, being minimal at the channel center and increasing near the walls. Thus, the morphological evolution of the droplets in the middle of the channel is practically unaffected by the microstructure destruction, their morphology remains constant regardless of the initial $f_{0}$. As, we move away from the center, in the case of $f_{0}=1$, the initially high viscosity $\eta=\eta_{\infty}(1+\alpha)$ limits deformation and $\lambda_{\text{thix}}$ governs the droplet response, with modest droplet deformation across the domain. In contrast, droplets with $f_{0}=0$, starting with the base viscosity $\eta_{\infty}$, tend to deform more strongly under flow leading to significantly more elongated droplets near the walls. The behavior observed in the emulsion of thixotropic droplets in a Newtonian continuous phase can be interpreted similarly to the previously discussed case. As in the previous scenario the thixotropic relaxation governs the evolution of the microstructure. Consistently, droplets traveling in the center of the channel retain their morphology regardless of the initial state $f_{0}$, whereas those closer to the walls experience more pronounced deformation due to larger Ca numbers. We must note that in the steady state, both systems ($f_{0}=0$ and $f_{0}=1$) should reach an equivalent state of droplets deformation. However, their transient evolution evidence the interplay between flow-induced stress and microstructural response, and highlight potential emergence of heterogeneous droplets morphology.
@@
Figure 8: Emulsion with Newtonian continuous phase and thixotropic dispersed phase. Comparison starting with the formed microstructure (top row) versus the unformed microstructure (bottom row). Here, $t^{*}=t/\lambda_{thix}$. The detail of the droplets is shown at $t^{*}=0.5$ (before reaching a steady state) to illustrate the differences in the transient state.